# CaveAgent: Transforming LLMs into Stateful Runtime Operators

### v2

@@ Abstract:
LLM-based agents are increasingly capable of complex task execution, yet current agentic systems remain constrained by text-centric paradigms that struggle with long-horizon tasks due to fragile multi-turn dependencies and context drift. We present CaveAgent, a framework that shifts tool use from “LLM-as-Text-Generator” to “LLM-as-Runtime-Operator.” CaveAgent introduces a dual-stream architecture that inverts the conventional paradigm: rather than treating the LLM’s text context as the primary workspace with tools as auxiliary, CaveAgent elevates the persistent Python runtime as the central locus of state, with a lightweight semantic stream serving as its orchestrator. Beyond leveraging code generation to resolve interdependent sub-tasks (e.g., loops, conditionals) in a single step, CaveAgent introduces Stateful Runtime Management: it injects, manipulates, and retrieves complex Python objects (e.g., DataFrames, database connections) that persist across turns, unlike existing code-based approaches that remain text-bound. CaveAgent further provides a runtime-integrated skill management system that extends the Agent Skills open standard, enabling ecosystem interoperability through executable skill injections. This persistence mechanism serves as a high-fidelity external memory that reduces context drift in multi-turn interactions and preserves processed data for downstream applications without information loss. Evaluations on Tau2-bench, BFCL, and case studies across representative SOTA LLMs demonstrate consistent improvements: a 10.5% success rate gain on retail tasks, 28.4% reduction in total token consumption for multi-turn scenarios, and 59% token reduction on data-intensive tasks, enabling CaveAgent to handle data scales that cause context overflow in both JSON-based and code-based agents. The accessible runtime state further provides programmatically verifiable feedback, enabling automated evaluation and reward signal generation without human annotation and establishing a structural foundation for future research in Reinforcement Learning with Verifiable Rewards (RLVR).
@@ 1 Introduction
We aim to build a system that utilizes Python’s "everything is an object” philosophy to enable full Object-Oriented function calling and interaction, delegating context engineering to a persistent runtime and allowing the direct injection and retrieval of high-fidelity objects without serialization loss, thereby fully leveraging the strong code-generation capability of LLMs.
@@
  • We introduce CaveAgent, a tool-use framework built on Stateful Runtime Management. CaveAgent shifts the paradigm from process-oriented function calling to persistent, object-oriented state management. It achieves context compression and context-grounded memory recall by delegating context engineering to a persistent runtime, eliminating the token overhead and precision loss of textual serialization while enabling few-step resolution of logically interdependent tasks. CaveAgent also provides runtime-integrated skill management that extends the Agent Skills open standard with executable injections for ecosystem interoperability.
@@ 2.1.3 Security Check via Static Analysis
2.1.4 Runtime-Integrated Skill Management
@@
CaveAgent extends the Agent Skills open standard [^2] https://agentskills.io, a portable specification adopted across the AI ecosystem including Claude Code, Gemini CLI, and Cursor. by introducing an injection.py module alongside the standard SKILL.md file. While standard skills provide text-based prompts that guide LLM behavior, CaveAgent skills additionally export Functions, Variables, and Type definitions that are injected directly into the persistent runtime upon activation. The framework employs progressive disclosure: skill metadata (name and description) is loaded at startup for routing decisions, while full instructions and runtime injections are loaded on-demand when the agent invokes activate_skill(). This bridges declarative skill definitions with CaveAgent’s stateful runtime paradigm: skills become executable capabilities, not merely textual instructions. As shown in Figure 4, domain expertise is packaged as both human-readable instructions for the language model and machine-executable artifacts for the runtime.
@@ 4 Conclusion
We present CaveAgent, a framework that transforms LLM tool use from stateless JSON function calling to persistent, object-oriented stateful runtime management. CaveAgent enables agents to maintain high-fidelity memory of complex objects and execute sophisticated logic via Python code. Experiments on Tau2-bench show that this approach outperforms SOTA baselines in multi-turn success rates (+10.5%) and token efficiency. Beyond performance gains, a key contribution is the programmatic verifiability enabled by CaveAgent’s architecture: because runtime state is deterministically inspectable, agent behavior can be evaluated automatically without human annotation, establishing a structural foundation for Reinforcement Learning with Verifiable Rewards and runtime-mediated multi-agent coordination. Qualitative case studies are provided in Appendix G.

### v3

@@ Abstract:
LLM-based agents are increasingly capable of complex task execution, yet current agentic systems remain constrained by text-centric paradigms that struggle with long-horizon tasks due to fragile multi-turn dependencies and context drift. We present CaveAgent, a framework that shifts LLM tool use from “LLM-as-Text-Generator” to “LLM-as-Runtime-Operator.” CaveAgent introduces a dual-stream architecture: a semantic stream for lightweight reasoning and a runtime stream backed by a persistent Python environment for stateful execution. Rather than treating the LLM’s text context as the primary workspace, CaveAgent elevates the persistent runtime as the central locus. Beyond leveraging code generation to resolve interdependent sub-tasks (e.g., loops, conditionals) in a single step, CaveAgent introduces Stateful Runtime Management: it injects, manipulates, and retrieves complex Python objects (e.g., DataFrames, database connections) that persist across turns, unlike existing code-based approaches that remain text-bound. CaveAgent further provides a runtime-integrated skill management system that extends the Agent Skills open standard, enabling ecosystem interoperability through executable skill injections. This persistence mechanism serves as a high-fidelity external memory that reduces context drift in multi-turn interactions and preserves processed data for downstream applications with less information loss. Evaluations on Tau2-bench and BFCL across six SOTA LLMs demonstrate consistent improvements: +10.5% success rate on retail tasks, 28.4% reduction in total token consumption, and 59% token reduction on data-intensive tasks. The accessible runtime state further provides programmatically verifiable feedback, enabling automated evaluation and reward signal generation without human annotation and establishing a structural foundation for future research in Reinforcement Learning with Verifiable Rewards (RLVR).
@@ 1 Introduction
We aim to build a system that utilizes Python’s "everything is an object” philosophy to enable full Object-Oriented function calling and interaction, delegating context engineering to a persistent runtime and allowing the direct injection and retrieval of high-fidelity objects without serialization loss, thereby fully leveraging the strong code-generation capability of LLMs, and enabling modular distribution of executable tool capabilities through a portable Agent-Skill standard.
@@
  • We introduce CaveAgent, a tool-use framework built on Stateful Runtime Management. CaveAgent shifts the paradigm from process-oriented function calling to persistent, object-oriented state management. It achieves context compression and context-grounded memory recall by delegating context engineering to a persistent runtime, eliminating the token overhead and precision loss of textual serialization while enabling few-step resolution of logically interdependent tasks. CaveAgent also introduces runtime-integrated skill management that extends the Agent Skills open standard: because the architecture treats tools as first-class Python objects, skills can deliver executable artifacts (functions, variables, and type definitions) directly into the runtime upon activation, unifying tool registration and skill distribution into a single mechanism.
@@ 2.1.3 Security Check via Static Analysis
2.2 Runtime-Integrated Skill Management
@@
CaveAgent extends the Agent Skills open standard [^2] https://agentskills.io, a portable specification adopted across the AI ecosystem including Claude Code, Gemini CLI, and Cursor. by introducing an injection.py module alongside the standard SKILL.md file. While standard skills provide text-based prompts that guide LLM behavior, CaveAgent skills additionally export Functions, Variables, and Type definitions that are injected directly into the persistent runtime upon activation. The framework employs progressive disclosure: skill metadata (name and description) is loaded at startup for routing decisions, while full instructions and runtime injections are loaded on-demand when the agent invokes activate_skill(). This bridges declarative skill definitions with CaveAgent’s stateful runtime paradigm: skills deliver executable artifacts into the runtime, not merely textual instructions to the LLM. As shown in Figure 4, domain expertise is packaged as both human-readable instructions for the language model and machine-executable artifacts for the runtime.
@@
Tool delivery Cannot deliver tools Injects tools into runtime
Discussion: From Tool Disclosure to Object Disclosure.
The skill activation pattern (lazy, metadata-driven retrieval of executable artifacts) generalizes beyond tools to any runtime object. Intermediate results stored as persistent variables (“Code as Action, State as Memory”) can similarly be discovered via semantic search over their metadata (names, types, descriptions), bringing relevant state into the agent’s context only when needed. Unlike retrieval-augmented generation over text chunks, the retrieved items are live Python objects with full type fidelity. This extends progressive disclosure from tool management to general workflow-state management.
@@ 4 Conclusion
We present CaveAgent, a framework that transforms LLM tool use from stateless JSON function calling to persistent, object-oriented stateful runtime management. CaveAgent enables agents to maintain high-fidelity memory of complex objects and execute sophisticated logic via Python code. By extending the Agent Skills open standard with runtime injection, CaveAgent demonstrates that the persistent runtime paradigm enables a new mode of tool distribution, skills that deliver executable artifacts rather than solely textual instructions, unifying tool registration and skill management into a single portable mechanism. Experiments on Tau2-bench show that this approach outperforms SOTA baselines in multi-turn success rates (+10.5%) and token efficiency. Beyond performance gains, a key contribution is the programmatic verifiability enabled by CaveAgent’s architecture: because runtime state is deterministically inspectable, agent behavior can be evaluated automatically without human annotation, establishing a structural foundation for Reinforcement Learning with Verifiable Rewards and runtime-mediated multi-agent coordination. Qualitative case studies are provided in Appendix G.